# main_arXiv.bib

### v1

@@ 3.1 Analysis of MB-RS-QVI
The overall proof technique for the upper bounds can also be used to derive upper bounds in the risk-neutral case by using the classical simulation lemma [kearns2002near]. Doing this will however yield a suboptimal upper bound. State-of-the-art techniques for proving optimal upper bounds in the risk-neutral case are however critically exploiting linearity of the expectation operator and are thus not readily available in the risk-sensitive case due to the non-linearity of the entropic risk measure.
@@ 4 Sample Complexity Lower Bounds
In this section, we provide two sample complexity lower bounds. The first one, presented in Theorem 3, concerns the sample complexity of learning the optimal $Q$-value function $Q^{*}$, whereas the second one, in Theorem 4, is on learning an optimal policy $\pi^{*}$. The proofs are both postponed to Appendix D.
@@
There exist constants $c_{1},c_{2}>0$ such that for any RL algorithm $\mathcal{U}$ that outputs a $Q$-function $Q^{\mathcal{U}}$ and any $\delta\in(0,\frac{1}{4})$ and $\varepsilon\in(0,\frac{1}{40}\frac{\gamma}{|\beta|}(1-e^{-|\beta|\frac{1}{1-\gamma}})$, the following holds: if the total number $T$ of transitions satisfies
$T\leq\frac{SA\gamma^{2}}{c_{1}\varepsilon^{2}}\frac{(e^{|\beta|\frac{1}{1-\gamma}}-3)}{|\beta|^{2}}\log\bigg{(}\frac{SA}{c_{2}\delta}\bigg{)}\,,$
@@
There exist constants $c_{1},c_{2}>0$ such that for any RL algorithm $\mathcal{U}$ that outputs a policy $Q^{\mathcal{U}}$ and any $\delta\in(0,\frac{1}{4})$ and $\varepsilon\in(0,\frac{1}{50}\frac{\gamma}{|\beta|}(1-e^{-|\beta|\frac{1}{1-\gamma}})$, it holds that if the total number $T$ of transitions satisfies
$T\leq\frac{SA\gamma^{2}}{c_{1}\varepsilon^{2}}\frac{(e^{|\beta|\frac{1}{1-\gamma}}-3)}{|\beta|^{2}}\log\bigg{(}\frac{S}{c_{2}\delta}\bigg{)},$
@@
The lower bounds in Theorems 3-4 establish that an exponential dependence of the sample complexity on the effective horizon $\frac{1}{1-\gamma}$; more precisely, they assert that a scaling of $e^{|\beta|\frac{1}{1-\gamma}}$ is unavoidable in both $Q$-learning and policy-learning. Recalling the sample complexity bounds of MB-RS-QVI (Theorems 1-2), we observe a similar exponential dependence. However, there remains a gap of order $\frac{1}{(1-\gamma)^{2}}e^{|\beta|\frac{1}{1-\gamma}}$. This remains as an interesting open question as to whether closing this gap can be done via a more elegant analysis of MB-RS-QVI or it calls for more novel algorithmic ideas.
The proofs of lower bounds are given in Section D in the appendix, but we briefly sketch below the ideas here for the case where the algorithm outputs a $Q$-function; the other case is proven using quite similar ideas.
@@
We consider a class of hard-to-learn MDPs, admitting a structure similar to that used in the lower bound derivation of [azar2011reinforcement]. In such MDPs, the state-action pairs are organized into triplets $\{z^{0},z^{G},z^{B}\}$, where $z^{G}$ yields a reward of $1$, $z^{B}$ yields a reward of 00, and both $z^{G}$ and $z^{B}$ are absorbing. In $z_{0}$, the reward is also zero and the transition function satisfies $P(z^{G}|z^{0})=q$ and $P(z^{B}|z^{0})=1-q$, for some $q>0$. This construction critically allows us to calculate explicitly $Q^{*}(z)$ for a given parameter $q$ and for two different MDPs $M_{0},M_{1}$ in the class where $q_{0}=p$ and $q_{1}=p+\alpha$ for appropriately chosen values of $p$ and $\alpha$ we are able to ensure that $Q^{*}_{M_{1}}(z)-Q^{*}_{M_{0}}(z)>2\varepsilon$ which means that any specific algorithmic output $Q^{\mathcal{U}}(z)$ cannot be $\varepsilon$-close to both $Q^{*}_{M_{1}}(z)$ and $Q^{*}_{M_{0}}(z)$. We then show by a likelihood ratio argument that any algorithm $\mathcal{U}$ that is $(\varepsilon,\delta)$-correct on $M_{0}$, i.e. that $\mathbb{P}_{0}(|Q_{M_{0}}^{*}(z)-Q^{\mathcal{U}}(z)|\leq\varepsilon)>\delta$, will also satisfy that $\mathbb{P}_{1}(|Q_{M_{0}}^{*}(z)-Q^{\mathcal{U}}(z)|\leq\varepsilon)>\delta$ provided that the algorithm does not try out $z$ enough times on $M_{0}$ and exactly because $Q^{*}_{M_{1}}(z)-Q^{*}_{M_{0}}(z)>2\varepsilon$, the event $\{|Q_{M_{0}}^{*}(z)-Q^{\mathcal{U}}(z)|\leq\varepsilon\}$ is disjoint from the event on being $\varepsilon$-close to $Q_{M_{1}}^{*}$. The final part of the proof is to exploit that the different triplets of state-action pairs contain no information about each other which allows for an independence argument for the estimation of $Q^{\mathcal{U}}(z)$ and $Q^{\mathcal{U}}(z^{\prime})$ for $z\neq z^{\prime}$. While doing this analysis, we fix an inaccuracy in the proof of Lemma 17 in [gheshlaghi2013minimax] that arises where they lower-bound the likelihood ratio of two Bernoulli random variables with biases $p\geq\frac{1}{2}$ and $p+\alpha$ on a high probability event. We also mention that we extend the result to hold for $p<\frac{1}{2}$.
@@
It is worth remarking that the best lower bound in the risk-neutral setting is derived in [gheshlaghi2013minimax] using a richer construction than above. However, with a risk-sensitive learning objective, the optimal state-action value function in the construction of [gheshlaghi2013minimax] does not admit an analytical solution, which is needed for the delicate tuning of the transition probabilities.
@@ 5 Conclusion and Future Works
We have studied the sample-complexity of learning the optimal $Q$-function and that of learning an optimal policy in finite discounted MDPs, where the agent has recursive risk-preferences given by the entropic risk measure and has access to a generative model. We introduced an algorithm, called MB-RS-QVI, and derived PAC-type bounds on its sample complexity for both learning which have derived bounds from analyzing the MB-RS-QVI  algorithm that uses the model given by the plug-in estimator from samples generated by a simulator. We also derive lower bounds that show that dependence on $e^{|\beta|\frac{1}{1-\gamma}}$ is unavoidable in both cases. The bounds that we derive on the sample complexity of learning the optimal $Q$-value are of order
$\mathcal{O}\bigg{(}\big{(}S+\log(SA/\delta)\big{)}\frac{SA}{\varepsilon^{2}(1-\gamma)^{2}}\frac{e^{2|\beta|\frac{1}{1-\gamma}}}{|\beta|^{2}}\bigg{)},\qquad\Omega\bigg{(}\frac{SA}{\varepsilon^{2}}\frac{e^{|\beta|\frac{1}{1-\gamma}}}{|\beta|^{2}}\log(SA/\delta)\bigg{)}$ (18)
@@
$\mathcal{O}\bigg{(}\big{(}S+\log(SA/\delta)\big{)}\frac{SA}{\varepsilon^{2}(1-\gamma)^{2}}\frac{e^{2|\beta|\frac{1}{1-\gamma}}}{|\beta|^{2}}\bigg{)},\qquad\Omega\bigg{(}\frac{SA}{\varepsilon^{2}}\frac{e^{|\beta|\frac{1}{1-\gamma}}}{|\beta|^{2}}\log(S/\delta)\bigg{)}$ (19)
These constitute the first bounds, to our knowledge, on the sample complexities of entropic risk-sensitive agents in the discounted MDP setting. The upper and lower bounds derived in this paper leave open gaps of sizes, at most, $\frac{S}{(1-\gamma)^{2}}e^{|\beta|\frac{1}{1-\gamma}}$. Since the constructions in the lower bounds are not the ones used in the tightest lower bounds derived in the risk-neutral setting, one possibility is that the lower bounds can be improved by considering a more carefully chosen set of hard-to-learn MDPs with the challenge being to control the gap in $V$-values or $Q$-values under different parameters. Also since the plugin-estimator model-based QVI algorithm is provably optimal in the risk-neutral setting, we believe that this might also be the case for risk-sensitive agents but that more tools are needed to develop a more careful analysis. Another future direction is that of developing model-free algorithms for this setting and analyzing their statistical efficiency. Another interesting research direction is to consider function approximation, as in [zhou2021provably]. As other future directions, one may consider more complicated RL settings such as offline RL [rashidinejad2021bridging], where data is collected under a fixed (but unknown) behavior policy, and online RL [strehl_analysis_2008, lattimore_near-optimal_2014], where the agent’s learned policy impacts the data collection process. And finally one may also consider the problem where the learner have non-recursive risk-preferences instead.
@@ Appendix A Technical Lemmas
The first lemma establishes a decomposition result for MB-RS-QVI, whose proof follows very similar lines to the proof of Lemma 3 in [agarwal_model-based_2020].
@@
$V^{\pi_{k}}(s)-V^{*}(s)$ $=V^{\pi_{k}}(s)-\hat{V}^{\pi_{k}}(s)+\hat{V}^{\pi_{k}}(s)-\hat{V}^{*}(s)+\hat{V}^{*}(s)-V^{*}(s)$
$\geq V^{\pi_{k}}(s)-\hat{V}^{\pi_{k}}(s)+\hat{V}^{\pi_{k}}(s)-\hat{V}^{*}(s)+\hat{V}^{\pi*}(s)-V^{*}(s)$
$\geq-\|V^{\pi_{k}}-\hat{V}^{\pi_{k}}\|-\|\hat{V}^{\pi_{k}}-\hat{V}^{*}\|-\|\hat{V}^{\pi^{*}}-V^{*}\|\,,$
@@
Let $f(x)=(1-x)$ and $g(x)=-x-x^{2}+x^{3}$. It holds that $f(0)=g(0)$, and since we have $f^{\prime}(x)=\frac{1}{1-x}$ and $g^{\prime}(x)=-1-2x+3x^{2}$, it follows easily that

### v2

@@ 3.1 Analysis of MB-RS-QVI
State-of-the-art techniques for proving optimal upper bounds in the risk-neutral case are critically exploiting linearity of the expectation operator and are thus not readily available in the risk-sensitive case due to the non-linearity of the entropic risk measure.
@@ 4 Sample Complexity Lower Bounds
In this section, we provide two sample complexity lower bounds. The first one, presented in Theorem 4, concerns the sample complexity of learning the optimal $Q$-value function $Q^{*}$, whereas the second one, in Theorem 5, is on learning an optimal policy $\pi^{*}$. The proofs are both postponed to Appendix H.
@@
There exist constants $c_{1},c_{2}>0$ such that for any RL algorithm $\mathcal{U}$ that outputs a $Q$-function $Q^{\mathcal{U}}$ and any $\delta\in(0,\frac{1}{4})$ and $\varepsilon\in\big(0,\frac{1}{40}\frac{\gamma}{|\beta|}(1-e^{-|\beta|\frac{1}{1-\gamma}})\big)$, the following holds: if the total number $T$ of transitions satisfies
$T\leq\frac{(S-2)A\gamma^{2}}{c_{1}\varepsilon^{2}}\frac{(e^{|\beta|\frac{1}{1-\gamma}}-3)}{|\beta|^{2}}\log\bigg(\frac{(S-2)A}{c_{2}\delta}\bigg)\,$,
@@
There exist constants $c_{1},c_{2}>0$ such that for any RL algorithm $\mathcal{U}$ that outputs a policy $Q^{\mathcal{U}}$ and any $\delta\in(0,\frac{1}{4})$ and $\varepsilon\in\big(0,\frac{1}{50}\frac{\gamma}{|\beta|}(1-e^{-|\beta|\frac{1}{1-\gamma}})\big)$, it holds that if the total number $T$ of transitions satisfies
$T\leq\frac{(S-2)(A-1)\gamma^{2}}{c_{1}\varepsilon^{2}}\frac{(e^{|\beta|\frac{1}{1-\gamma}}-3)}{|\beta|^{2}}\log\bigg(\frac{(S-2)}{c_{2}\delta}\bigg),$
@@
The lower bounds in Theorems 4-5 establish that an exponential dependence of the sample complexity on the effective horizon $\frac{1}{1-\gamma}$ is unavoidable; more precisely, they assert that a scaling of $e^{|\beta|\frac{1}{1-\gamma}}$ is unavoidable in both $Q$-learning and policy-learning. Recalling the sample complexity bounds of MB-RS-QVI (Theorems 1-2), we observe a similar exponential dependence. However, there remains a gap of order $\frac{1}{(1-\gamma)^{2}}e^{|\beta|\frac{1}{1-\gamma}}$. This remains as an interesting open question as to whether closing this gap can be done via a more elegant analysis of MB-RS-QVI or it calls for more novel algorithmic ideas, but in any case the lower bounds show that risk-sensitive agents face a fundamentally harder problem than risk-neutral agents where the sample complexity is polynomial in $\frac{1}{1-\gamma}$. The proofs of lower bounds are given in Section H in the appendix, but we briefly sketch the ideas below for the case where the algorithm outputs a $Q$-function; the other case is proven using quite similar ideas.
@@
We consider a class of hard-to-learn MDPs. In the class there are two absorbing states $s^{G}$ and $s^{B}$ where in $s^{G}$ the agent always recieves a reward of $R=1$ and in $s^{B}$ the agent always receives a reward of $R=0$ irrespective of the action taken. For all other state-action pairs $z$ the reward is zero and the only possible transition is to either of the states $s^{G}$ and $s^{B}$. $P(s^{G}|z)=q$ and $P(s^{B}|z)=1-q$, for some $q>0$. This construction critically allows us to calculate explicitly $Q^{*}(z)$ for a given parameter $q$ and for two different MDPs $M_{0},M_{1}$ in the class where $q_{0}=p$ and $q_{1}=p+\alpha$ for appropriately chosen values of $p$ and $\alpha$ we are able to ensure that $Q^{*}_{M_{1}}(z)-Q^{*}_{M_{0}}(z)>2\varepsilon$ which means that any specific algorithmic output $Q^{\mathcal{U}}(z)$ cannot be $\varepsilon$-close to both $Q^{*}_{M_{1}}(z)$ and $Q^{*}_{M_{0}}(z)$. We then show by a likelihood ratio argument that any algorithm $\mathcal{U}$ that is $(\varepsilon,\delta)$-correct on $M_{0}$, i.e. that $\mathbb{P}_{0}(|Q_{M_{0}}^{*}(z)-Q^{\mathcal{U}}(z)|\leq\varepsilon)>\delta$, will also satisfy that $\mathbb{P}_{1}(|Q_{M_{0}}^{*}(z)-Q^{\mathcal{U}}(z)|\leq\varepsilon)>\delta$ provided that the algorithm does not try out $z$ enough times on $M_{0}$ and exactly because $Q^{*}_{M_{1}}(z)-Q^{*}_{M_{0}}(z)>2\varepsilon$, the event $\{|Q_{M_{0}}^{*}(z)-Q^{\mathcal{U}}(z)|\leq\varepsilon\}$ is disjoint from the event on being $\varepsilon$-close to $Q_{M_{1}}^{*}$. The final part of the proof is to exploit that the different state-action pairs contain no information about each other which allows for an independence argument for the estimation of $Q^{\mathcal{U}}(z)$ and $Q^{\mathcal{U}}(z^{\prime})$ for $z\neq z^{\prime}$. While doing this analysis, we fix an inaccuracy in the proof of Lemma 17 in gheshlaghi2013minimax that arises where they lower-bound the likelihood ratio of two Bernoulli random variables with biases $p\geq\frac{1}{2}$ and $p+\alpha$ on a high probability event. We also mention that we extend the result to hold for $p<\frac{1}{2}$.
For policy learning we consider almost the same class of MDPs but augment with a known state $a_{0}$ that is used in the analysis.
@@
It is worth remarking that the best lower bound in the risk-neutral setting is derived in gheshlaghi2013minimax using a richer construction than above. However, with a risk-sensitive learning objective, the optimal state-action value function in the construction of gheshlaghi2013minimax does not admit an analytical solution, which is needed for the delicate tuning of the transition probabilities.
@@ 5 Conclusion and Future Works
We have studied the sample-complexity of learning the optimal $Q$-function and that of learning an optimal policy in finite discounted MDPs, where the agent has recursive risk-preferences given by the entropic risk measure and has access to a generative model. We introduced an algorithm, called MB-RS-QVI, and derived PAC-type bounds on its sample complexity for both learning which have derived bounds from analyzing the MB-RS-QVI algorithm that uses the model given by the plug-in estimator from samples generated by a simulator. We also derive lower bounds. The upper bounds show that PAC-learning is possible but the lower bounds show that dependence on $e^{|\beta|\frac{1}{1-\gamma}}$ is unavoidable and thus that learning is fundamentally harder for risk-senstive agents relative to risk-neutral agents. The bounds that we derive on the sample complexity of learning the optimal $Q$-value are of order
$\mathcal{O}\bigg(\log(SA/\delta)\big)\frac{SA}{\varepsilon^{2}(1-\gamma)^{2}}\bigg(\frac{e^{|\beta|\frac{1}{1-\gamma}}-1}{|\beta|}\bigg)^{2}\bigg),\qquad\Omega\bigg(\frac{(S-2)A}{\varepsilon^{2}}\frac{e^{|\beta|\frac{1}{1-\gamma}}-3}{|\beta|^{2}}\log((S-2)A/\delta)\bigg)$ (17)
@@
$\mathcal{O}\bigg(\log(SA/\delta)\big)\frac{SA}{\varepsilon^{2}(1-\gamma)^{4}}\bigg(\frac{e^{|\beta|\frac{1}{1-\gamma}}-1}{|\beta|}\bigg)^{2}\bigg),\qquad\Omega\bigg(\frac{(S-2)(A-1)}{\varepsilon^{2}}\frac{e^{|\beta|\frac{1}{1-\gamma}}-3}{|\beta|^{2}}\log((S-2)/\delta)\bigg)$ (18)
where we also give the alternative bound of order
$\qquad\mathcal{O}\bigg(\big(S+\log(SA/\delta)\big)\frac{SA}{\varepsilon^{2}(1-\gamma)^{2}}\frac{e^{2|\beta|\frac{1}{1-\gamma}}}{|\beta|^{2}}\bigg)$ (19)
which might be tighter in cases of long horizon and small state space. These constitute the first bounds, to our knowledge, on the sample complexities of entropic risk-sensitive agents in the discounted MDP setting. The upper and lower bounds derived in this paper leave open gaps in $\frac{1}{1-\gamma}$. Since the constructions in the lower bounds are not the ones used in the tightest lower bounds derived in the risk-neutral setting, one possibility is that the lower bounds can be improved by considering a more carefully chosen set of hard-to-learn MDPs with the challenge being to control the gap in $V$-values or $Q$-values under different parameters. Also since the plugin-estimator model-based QVI algorithm is provably optimal in the risk-neutral setting, we believe that this might also be the case for risk-sensitive agents but that more tools are needed to develop a more careful analysis. Another future direction is that of developing model-free algorithms for this setting and analyzing their statistical efficiency. Another interesting research direction is to consider function approximation, as in zhou2021provably . As other future directions, one may consider more complicated RL settings such as offline RL rashidinejad2021bridging , where data is collected under a fixed (but unknown) behavior policy, and online RL strehl_analysis_2008 ; lattimore_near-optimal_2014 , where the agent’s learned policy impacts the data collection process. And finally one may also consider the problem where the learner have non-recursive risk-preferences instead.
@@ Appendix A Technical Lemmas
The first lemma establishes a decomposition result for MB-RS-QVI, whose proof follows very similar lines to the proof of Lemma 3 in [24].
@@
$V^{\pi_{k}}(s)-V^{*}(s)$ $=V^{\pi_{k}}(s)-\widehat{V}^{\pi_{k}}(s)+\widehat{V}^{\pi_{k}}(s)-\widehat{V}^{*}(s)+\widehat{V}^{*}(s)-V^{*}(s)$
$\geq V^{\pi_{k}}(s)-\widehat{V}^{\pi_{k}}(s)+\widehat{V}^{\pi_{k}}(s)-\widehat{V}^{*}(s)+\widehat{V}^{\pi*}(s)-V^{*}(s)$
$\geq-\|V^{\pi_{k}}-\widehat{V}^{\pi_{k}}\|-\|\widehat{V}^{\pi_{k}}-\widehat{V}^{*}\|-\|\widehat{V}^{\pi^{*}}-V^{*}\|\,$,
@@
Let f(x)=(1x)f(x)=(1-x) and $g(x)=-x-x^{2}+x^{3}$. It holds that $f(0)=g(0)$, and since we have $f^{\prime}(x)=\frac{1}{1-x}$ and $g^{\prime}(x)=-1-2x+3x^{2}$, it follows easily that